# Data for Mathematical Copilots: Better Ways of Presenting Proofs for Machine Learning

### v1

@@ 3 The Past – Common Pitfalls in Current Datasets
In this section, we discuss several issues with current datasets. Note that our limitations are with respect to the landscape of AI for mathematics at the time of writing (Dec 2024); it is possible that some of these limitations are addressed. Indeed, we hope that they are addressed swiftly. We first list issues that are common to both natural-language datasets and datasets comprising formalized mathematics. Then, we list issues that are specific to either natural-language datasets or formal datasets.
@@ 3.1.1 Difficulty
Several datasets have been proposed which explore proof sophistication, such as the OlympiadBench (He et al., 2024), the IMO Small Challenge, or the IMO-AG-30 dataset (Trinh et al., 2024), but very few existing datasets reach a level of difficulty that is more advanced than that of an undergraduate degree in mathematics in terms of level of abstraction. For a (non-exhaustive) overview of datasets, ordered by their difficulty, see Table 1.
For example, MATH focuses on high-school competition problems, while the recent TheoremQA (Chen et al., 2023) and Advanced Reasoning Benchmark (ARB) (Sawada et al., 2023) are upper-undergraduate level. Many commonly used datasets test mathematics below the undergraduate level, such as GSM8k and the 23 tasks in the Lila benchmark (Mishra et al., 2022b).
Two datasets that go beyond upper-undergraduate level are the GHOSTS dataset (Frieder et al., 2023c), as well as the NaturalProofs(-Gen) (Welleck et al., 2021, 2022) datasets, which contain problems, theorems, and proofs on a wide range of more advanced mathematical topics. However, these suffer from a lack of automated evaluation and contain many problems that could, in principle, be in the training corpora of modern LLMs, as we discuss later. FrontierMath (Glazer et al., 2024) is also a more advanced dataset, but is currently not publicly accessible.
Most current LLMS (e.g., the Qwen (Yang et al., 2024a) family of models, or commercial models released by OpenAI and Google) are evaluated on the MATH dataset (Hendrycks et al., 2021) and the GSM8K dataset (Cobbe et al., 2021), with a recent trend of some LLMs using held-out datasets such as the Hungarian National Finals high-school exam as an additional test of generalization (Paster, 2023). The MATH dataset uses five categories of difficulty and is strictly more difficult than GSM8K. Nonetheless, state-of-the-art LLMs have been found to reach an acceptable performance on it (Frieder et al., 2023a). For instance, Gemini reaches a performance of about 50%. [^5] https://blog.google/technology/ai/google-gemini-ai/performance, while solving almost all problems on GSM8K. Progress in LLMs is proceeding at a rapid pace, and the “Math-Specialized” version of Gemini 1.5 reaches (at best) a score of 91% on MATH, which is representative of the current state of the art (although, due to the closed nature of this line of research, it is difficult to place this achievement into context, such as by analyzing how similar training data has been to the MATH dataset, and how much was needed).
@@
This is particularly striking in the case of AlphaGeometry, which can solve 25 geometry problems from the IMO out of a test set of 30 problems (Trinh et al., 2024) (and 213 out of 231 on a larger benchmark). We argue that the accompanying dataset, while indeed difficult in terms of proof sophistication, is not sufficiently large to be able to accurately compare AlphaGeometry with other models.
@@
We believe that it is crucial to introduce more difficult datasets into the evaluation repertoire of math-AI. And importantly, the skill level of these datasets should be graded, such that we can form a nuanced characterization of model capabilities – this can be achieved, in part, by interpolating difficulty level between existing datasets and those representative of advanced mathematics. This includes datasets on graduate-level mathematical domains, such as functional analysis (to measure handling of arguments involving high levels of abstraction), partial differential equations (to measure handling of intricate computational arguments), as well as other domains that come with their own “mathematical flavor”. At the same time, more comprehensive datasets illustrating and measuring advanced problem-solving skills are needed. The miniF2F (Zheng et al., 2022) dataset and the GHOSTS dataset are just the first steps in this direction. For natural language datasets, we note that simply scraping many competition-level problem sets, coupled with an evaluation principle that only verifies a “final answer”, is not sufficient to ensure that systems can progress to becoming useable mathematical copilots. Such problem sets, even if some of them ultimately are only math word problems, require expert, specialized knowledge and training for humans to be proficient. Without a detailed evaluation protocol (see Section 3.1.2), not much insight will be gained into why poor performance was obtained - and, conversely, human evaluation will be very costly (see Section 3.2.1).
@@ 3.1.2 Binary Benchmarks
GHOSTS and NaturalProofs-Gen address this problem by using labels that are attached to language model outputs, which provide a classification of (mathematical) errors. For instance, GHOSTS distinguishes two types of labels, called “error codes” and “warning codes”, which make up, in total, 15 error or warning labels, while NaturalProofs-Gen distinguishes between 11 types of errors.
A small number of datasets use non-binary error rating schemes, to our knowledge these are only TheoremQA (3-point scheme) and ARB (5-point).
@@
We believe that it is crucial to develop datasets that support multi-aspect feedback. Relying on a single numerical representation, in the face of the diversity of mathematical failure modes, risks providing a signal that is too weak to be suitable for learning or evaluation – which is especially important to guard against when we consider deploying systems alongside people, who may care not only that a model is correct but also that the response is appropriately helpful (Collins et al., 2024a).
One of the reasons that AI researchers may have avoided multi-aspect feedback, particularly subjective feedback, is for fear of the difficulty of obtaining consistent human annotations. While annotator disagreement is a challenge, we suggest that the math-AI community look to the ways in which computer vision researchers have worked on developing new theoretical and empirical tools which lean in to the diversity of human responses (Uma et al., 2021; Collins et al., 2022; Sucholutsky et al., 2023; Gordon et al., 2022). Appropriately interleaving formality and subjectivity – especially for advanced mathematics, where there is a limited pool of possible human annotators – is a challenge, but one which demands cross-disciplinary collaboration and is ripe for future work.
As a middle-ground, we also recommend considering evaluation methods that do not assign numerical ratings at all – but rather annotate the outputs of AI systems with error codes and warning codes, then derive ratings from the codes directly (where no error codes mean a perfect rating).
@@ 3.1.3 Standardized Interaction Modes
Almost all datasets to date that are designed to evaluate LLMs capture only textbook-like questions. Moreover, owing to issues of automatic evaluation described in Section 3.2.1, the majority of datasets focus on a smaller subset of questions with a unique, typically numerical, answer. While such questions might span a range of different topics, from numerical (e.g., GSM8K, problems of level 1-2 from MATH, MultArith) and symbolic evaluations (the dataset associated with the work of (Lample and Charton, 2020)) to somewhat complex problem-solving tasks (problems of level 4-5 from MATH), they do not come close to covering all aspects of the mathematical tasks a mathematician might encounter in their daily practice. We believe that any future mathematical assistant should provide support for these tasks.
@@
    Proof-Specific Questions as the concept of a mathematical proof is varied and gives rise to a host of specific questions about proofs, which go beyond the type of questions related to proofs from the previous point. These are advanced questions concerning matters such as: Establishing whether a proof is effective or not, in the sense of proving a general statement about a mathematical object merely by proving its existence as opposed to producing a witness or an algorithm that computes; investigating what distinct proof strategies might be used to prove a certain statement; exploring the distinction between whether an “elementary” or an “advanced” proof is given.
@@
    Mathematical explanations and intuition, e.g., by providing (counter-)examples. Such questions are open-ended and have potentially (infinitely) many correct answers. This freeform exploration is crucial in the process of establishing the truth value of a given (advanced) mathematical statement, where one swings between attempting a direct proof of the statement and finding a counterexample to the statement.
@@
    Retrieval Tasks for definitions and mathematical facts, either from name to statement or reverse (“mathematical search engine”). This can range from a well-defined answer (e.g., Define this mathematical object!) to open-ended questions (what are the most important theorems in a certain field?; which results are relevant to prove a certain statement?; has a form of this statement already been proved?). This tests the capabilities of mathematical assistants to provide high-level explanations similar to those of a human expert in a certain mathematical field.
@@
    Informal Proof Completion, where a proof (or generally a mathematical statement) needs to be checked or completed. In the best case, the mathematical assistant should be able to provide a correction and fill in gaps. In such functionality, mathematical assistants can be running in the background to act as “co-pilots” and “mathematical grammar-checkers” — as the current generation of copilots [^6] Such as GitHub’s Copilot: https://github.com/features/copilot does successfully for programmers.
@@
    Mixed-mode interactions, where translation between formal and informal mathematics needs to be performed. These are particularly important for the combination of ITPs and general-purpose LLMs, which has recently picked up interest (e.g., (Jiang et al., 2023; First et al., 2023; Frieder et al., 2023c)).
@@
    Interactive, where the prompter (whose mathematical abilities can span a wide spectrum) solves mathematical questions together with the language model in an interactive way as in Collins et al. (2024a).
In practice, typical mathematical workflows require a mixture of different question and session types. Even if a question is not open-ended, the amount of mathematical detail and rigor the answer should contain will depend on the user’s background. In part, this should be deduced from the way the question is phrased. Currently, no datasets offer support for this.
We note that with the exception of Collins et al. (2024a) and the collection of CheckMate, almost all datasets are zero-shot. Interactive training modes are extensively employed in a formal language setup.
@@
  • Refactoring formal mathematics to yield nicer-looking proof is a significant undertaking. A tool that takes as input a formal proof and refactors it along desired criteria would be a welcome addition to the toolbox of researchers working in formalizing mathematics. To our knowledge, the only work to date that attempts this is the ImProver (Ahuja et al., 2024) for Lean, which allows formal proof creation of specific length.
@@ 3.1.4 Contamination
When releasing a dataset, control is lost over whether the dataset ends up as training data for machine learning models. For several of the state-of-the-art models, such as GPT-4 (OpenAI, 2023), that perform well on benchmarks, no information is available in the ensuing on whether efforts were made to ensure that training datasets were decontaminated and no datapoints from the evaluation benchmark were included in the training set. For Gemini 1.5 (Reid et al., 2024), Qwen2 (Yang et al., 2024a) and DeepSeekMath (Shao et al., 2024), only $n$-gram decontamination approaches have been tested. For complex datapoints, such as math, $n$-grams are unfortunately not always sufficient to ensure the training dataset is clean. There is evidence that suggests that several such modern models are already contaminated (Xu et al., 2024). While mitigating approaches have been proposed, such as generating new data, either from scratch (Mishra et al., 2024) or by using existing datasets as seeds (Zhou et al., 2024), it is not clear whether these can scale to more sophisticated mathematics, as the methods were tested using mathematics on the level of grade-school, and it is unclear how well they would scale to much higher levels of mathematical abstractions and problem-solving difficulty (the two main metrics of difficulty, as outlined in 3.1.1.
@@
If the dataset is sufficiently large, it is advisable to keep a part of the dataset hidden from the public so that the dataset creators can compare how well newly-released, publicly accessible open-weight models score in the public versus the hidden part of the dataset. Diverging scores on two splits of the dataset can indicate contamination. Care must be taken that datapoints from the hidden datasets are from the same distribution compared. For mathematics, this means at minimum: Same domain, difficulty range, and proof technique. This can be challenging: Given a specific problem $P$, it takes effort to source a new problem $P^{\star}$, that on all relevant metrics (including the mentioned ones) is similar to $P$, but at the same time is not completely analogous to $P$ – as it would be by, e.g., merely changing items in the problem statement that have a negligible on the proof (this is an approach followed by GSM8K-Symbolic dataset (Mirzadeh et al., 2024), which nonetheless turned out to be challenging for LLMs, highlighting their current limited reasoning capabilities). The approach advocated here has limitations for models released as-a-service (La Malfa et al., 2024), via APIs or GUIs, as running the hidden dataset through risks it automatically being used at a later stage as training data for those models.

### v2

@@ 3 The Past – Common Pitfalls in Current Datasets
In this section, we discuss several issues with current datasets. Note that our limitations are with respect to the landscape of AI for mathematics at the time of writing (Dec 2024); it is possible that some of these limitations are addressed by the time the reader reads this article. Indeed, we hope that they are addressed swiftly.
Underlying several of these problems is a dynamic described by Goodhart’s law: “When a measure becomes a target, it ceases to be a good measure.” A related formulation, Campbell’s law (campbell1979assessing), warns that quantitative indicators used for decision-making become increasingly subject to corruption pressures. In the context of mathematical AI, when LLM developers optimize for specific benchmark scores (such as accuracy on MATH or GSM8K), models may learn to exploit superficial patterns or benefit from data contamination rather than developing robust mathematical reasoning. High benchmark scores then no longer reliably indicate the capabilities that matter for mathematical copilots. This motivates our call for richer evaluation frameworks that are harder to “game” and that measure diverse facets of mathematical competence. Notably, the NLP community identified these same issues years ago and proposed solutions such as dynamic, adversarial benchmark creation (kiela-etal-2021-dynabench), yet the math-AI community has largely continued to rely on static benchmarks susceptible to the same failure modes.
Recent near-saturation results on boxed-answer benchmarks (deepseek2025r1; qwen2024qwq32bpreview) are consistent with this dynamic: as optimization concentrates on a small set of public targets, benchmark scores become less informative about broader mathematical competence.
We first list issues that are common to both natural-language datasets and datasets comprising formalized mathematics. Then, we list issues that are specific to either natural-language datasets or formal datasets.
@@ 3.1.1 Difficulty
Several datasets have been proposed which explore proof sophistication, such as the OlympiadBench (he2024olympiadbench), the IMO Small Challenge, or the IMO-AG-30 dataset (trinh2024solving), but very few existing datasets reach a level of difficulty that is more advanced than that of an undergraduate degree in mathematics in terms of level of abstraction. For a (non-exhaustive) overview of datasets, ordered by their difficulty, see Table 3.
For example, MATH focuses on high-school competition problems, while the recent TheoremQA (chen2023theoremqa) and Advanced Reasoning Benchmark (ARB) (sawada2023arb) are upper-undergraduate level. Many commonly used datasets test mathematics below the undergraduate level, such as GSM8k and the 23 tasks in the Lila benchmark (mishra-etal-2022-lila).
Two datasets that go beyond upper-undergraduate level are the GHOSTS dataset (frieder2023llms), as well as the NaturalProofs(-Gen) (welleck2021naturalproofs; welleck2022naturalprover) datasets, which contain problems, theorems, and proofs on a wide range of more advanced mathematical topics. However, these suffer from a lack of automated evaluation and contain many problems that could, in principle, be in the training corpora of modern LLMs, as we discuss later. FrontierMath (glazer2024frontiermath) is also a more advanced dataset, but is currently not publicly accessible. STORM-BORN is a recent derivation-style benchmark curated from recent research papers; it reports that strong models solve only a small fraction of the selected problems without task-specific training (stormborn2025).
Most current LLMS (e.g., the Qwen (yang2024qwen2) family of models, or commercial models released by OpenAI and Google) are evaluated on the MATH dataset (hendrycks2021measuring) and the GSM8K dataset (cobbe2021training), with a recent trend of some LLMs using held-out datasets such as the Hungarian National Finals high-school exam as an additional test of generalization (testing_language_models_on_a_held_out_high_school_national_finals_exam). The MATH dataset uses five categories of difficulty and is strictly more difficult than GSM8K. Nonetheless, state-of-the-art LLMs have been found to reach an acceptable performance on it (frieder2023mathematical). For instance, Gemini reaches a performance of about 50%. [^5] https://blog.google/technology/ai/google-gemini-ai/performance, while solving almost all problems on GSM8K. Progress in LLMs is proceeding at a rapid pace, and the “Math-Specialized” version of Gemini 1.5 reaches (at best) a score of 91% on MATH, which is representative of the current state of the art (although, due to the closed nature of this line of research, it is difficult to place this achievement into context, such as by analyzing how similar training data has been to the MATH dataset, and how much was needed).
@@
This is particularly striking in the case of AlphaGeometry, which can solve 25 geometry problems from the IMO out of a test set of 30 problems (trinh2024solving) (and 213 out of 231 on a larger benchmark). We argue that the accompanying dataset, while indeed difficult in terms of proof sophistication, is not sufficiently large to be able to accurately compare AlphaGeometry with other models.
Moreover, recent work casts doubt on whether high benchmark scores reflect genuine mathematical reasoning. nezhurina2024alice demonstrate that GPT-4, Claude 3 Opus, and other state-of-the-art LLMs at the time, exhibit dramatic reasoning failures on simple common-sense math problems (the “Alice in Wonderland” or AIW problems) that are trivially solvable by humans. These failures persist even with chain-of-thought prompting and multi-step re-evaluation, and models express strong overconfidence in their incorrect answers. If LLMs struggle with such elementary problems, it is difficult to credit their high performance on substantially harder benchmarks like MATH to genuine reasoning rather than to memorization or pattern-matching on familiar problem types. Proof or Bluff? reports expert grading of model solutions on the 2025 USAMO and finds low proof scores even for models that score well on short-answer benchmarks (prooforbluff2025). This tension motivates two directions: first, the development of out-of-distribution evaluation sets — for instance, via procedural generation of novel problems — that are less susceptible to memorization; and second, decontamination methods tailored specifically to mathematics, which must account for the fact that mathematically equivalent problems can appear syntactically different.
@@
We believe that it is crucial to introduce more difficult datasets into the evaluation repertoire of math-AI. And importantly, the skill level of these datasets should be graded, such that we can form a nuanced characterization of model capabilities – this can be achieved, in part, by interpolating difficulty level between existing datasets and those representative of advanced mathematics. This includes datasets on graduate-level mathematical domains, such as functional analysis (to measure handling of arguments involving high levels of abstraction), partial differential equations (to measure handling of intricate computational arguments), as well as other domains that come with their own “mathematical flavor”. At the same time, more comprehensive datasets illustrating and measuring advanced problem-solving skills are needed. The miniF2F (zheng2021minif2f) dataset and the GHOSTS dataset are just the first steps in this direction. For natural language datasets, we note that simply scraping many competition-level problem sets, coupled with an evaluation principle that only verifies a “final answer”, is not sufficient to ensure that systems can progress to becoming useable mathematical copilots. Such problem sets, even if some of them ultimately are only math word problems, require expert, specialized knowledge and training for humans to be proficient. Without a detailed evaluation protocol (see Section 3.1.2), not much insight will be gained into why poor performance was obtained - and, conversely, human evaluation will be very costly (see Section 3.2.1).
@@ 3.1.2 Binary Benchmarks
Benchmark diagnosticity and “training on the test task.”
A further limitation of many mathematics benchmarks is that they offer little insight into why a model succeeded or failed, beyond a single correctness bit. In practice, this can make it hard to improve a system other than by training on large quantities of data that is closely matched to the benchmark task distribution. Dominguez-Olmedo et al. (dominguez2024trainingonthetesttask) describe this as training on the test task: unlike leakage or contamination, it is a legitimate use of knowledge about benchmark tasks at training time, but it can confound both relative model comparisons and claims about emergent abilities. In mathematics, the problem is amplified by the concentration of benchmarks on a small set of task formats; Frieder et al. (frieder2023mathematical) describe a “vast unpopulated landscape of benchmarks” beyond the few overstudied datasets. One response is to make benchmarks more diagnostic by attaching structured feedback to failures, such as error codes (as in GHOSTS) that distinguish common mathematical failure modes, which can guide targeted data curation and model improvements rather than generic task-matching fine-tuning.
GHOSTS and NaturalProofs-Gen address this problem by using labels that are attached to language model outputs, which provide a classification of (mathematical) errors. For instance, GHOSTS distinguishes two types of labels, called “error codes” and “warning codes”, which make up, in total, 15 error or warning labels, while NaturalProofs-Gen distinguishes between 11 types of errors. A small number of datasets use non-binary error rating schemes, to our knowledge these are only TheoremQA (3-point scheme) and ARB (5-point).
@@
We believe that it is crucial to develop datasets that support multi-aspect feedback. Relying on a single numerical representation, in the face of the diversity of mathematical failure modes, risks providing a signal that is too weak to be suitable for learning or evaluation – which is especially important to guard against when we consider deploying systems alongside people, who may care not only that a model is correct but also that the response is appropriately helpful (collins2023evaluating).
One of the reasons that AI researchers may have avoided multi-aspect feedback, particularly subjective feedback, is for fear of the difficulty of obtaining consistent human annotations. While annotator disagreement is a challenge, we suggest that the math-AI community look to the ways in which computer vision researchers have worked on developing new theoretical and empirical tools which lean in to the diversity of human responses (uma2021learning; collins2022eliciting; sucholutsky2023informativeness; gordon2022jury). Appropriately interleaving formality and subjectivity – especially for advanced mathematics, where there is a limited pool of possible human annotators – is a challenge, but one which demands cross-disciplinary collaboration and is ripe for future work. As a middle-ground, we also recommend considering evaluation methods that do not assign numerical ratings at all – but rather annotate the outputs of AI systems with error codes and warning codes, then derive ratings from the codes directly (where no error codes mean a perfect rating).
@@ 3.1.3 Standardized Interaction Modes
Almost all datasets to date that are designed to evaluate LLMs capture only textbook-like questions. Moreover, owing to issues of automatic evaluation described in Section˜3.2.1, the majority of datasets focus on a smaller subset of questions with a unique, typically numerical, answer. While such questions might span a range of different topics, from numerical (e.g., GSM8K, problems of level 1-2 from MATH, MultArith) and symbolic evaluations (the dataset associated with the work of (lample2019deep)) to somewhat complex problem-solving tasks (problems of level 4-5 from MATH), they do not come close to covering all aspects of the mathematical tasks a mathematician might encounter in their daily practice. We believe that any future mathematical assistant should provide support for these tasks.
@@
    Proof-Specific Questions as the concept of a mathematical proof is varied and gives rise to a host of specific questions about proofs, which go beyond the type of questions related to proofs from the previous point. These are advanced questions concerning matters such as: Establishing whether a proof is effective or not, in the sense of proving a general statement about a mathematical object merely by proving its existence as opposed to producing a witness or an algorithm that computes; investigating what distinct proof strategies might be used to prove a certain statement; exploring the distinction between whether an “elementary” or an “advanced” proof is given. For example, the prime number theorem has classical complex-analytic proofs (via properties of the Riemann zeta function) (hadamard1896distribution) as well as later “elementary” proofs (selberg1949elementary; erdos1949newmethod).
@@
    Pictorial representation. Many mathematical objects and arguments have standard diagrammatic encodings that support reasoning and communication. Examples include plotting functions to inspect qualitative behavior (monotonicity, convexity, asymptotics), diagram-based proofs and explanations (for example, “proofs without words” collections (nelsen1993proofs)), and commutative or string diagrams used in areas such as algebra, topology, and category theory (selinger2011survey). A mathematical copilot that supports research practice should therefore be evaluated on tasks that translate between symbolic or textual descriptions and graphical representations, and on tasks that interpret diagrams to extract relevant properties or propose next steps in an argument.
@@
    Mathematical explanations and intuition, e.g., by providing (counter-)examples. Such questions are open-ended and have potentially (infinitely) many correct answers. This freeform exploration is crucial in the process of establishing the truth value of a given (advanced) mathematical statement, where one swings between attempting a direct proof of the statement and finding a counterexample to the statement. Terence Tao has recounted encountering a counterintuitive statement in compressed sensing, first attempting to construct a counterexample, and then redirecting that effort into a proof (bernstein2019mind).
    One recent counterexample-focused benchmark is CounterMATH, which asks models to refute undergraduate-level statements by producing explicit counterexamples (1,216 statement-rationale pairs) (countermath2025).
@@
    Proficiency in using mathematical tools, where the assistant is given access to external tools such as computer algebra systems (SageMath, Mathematica, Maple), numerical solvers, and domain-specific packages (for example, Singular, GAP, SnapPy), and is evaluated on four steps: selecting an appropriate tool, translating the mathematical intent into the tool’s input language, validating and interpreting the tool’s output, and composing multiple tool calls into a coherent mathematical argument. Tool-integrated reasoning approaches already use this pattern in restricted form (for example, delegating substeps to executable code) (gou2023tora). This also clarifies why Lean alone does not cover the full workflow of research mathematics: Lean can verify formal proofs, but many routine sub-tasks (symbolic simplification, numeric exploration, combinatorial search, heuristic experimentation) are more naturally handled by specialized systems. A practical copilot therefore needs a federation of tools, each with its own interface and notion of correctness, and datasets should record tool traces so that evaluation can replay the same sequence of calls.
@@
    Retrieval Tasks for definitions and mathematical facts, either from name to statement or reverse (“mathematical search engine”). MIRB provides a benchmark suite for mathematical retrieval, including premise retrieval and formula retrieval, which can serve as components of literature-search workflows (mirb2025). Retrieval tasks can range from a well-defined answer (e.g., Define this mathematical object!) to open-ended questions (what are the most important theorems in a certain field?; which results are relevant to prove a certain statement?; has a form of this statement already been proved?). This tests the capabilities of mathematical assistants to provide high-level explanations similar to those of a human expert in a certain mathematical field.
  7. 7.
    Informal Proof Completion, where a proof (or generally a mathematical statement) needs to be checked or completed. In the best case, the mathematical assistant should be able to provide a correction and fill in gaps. In such functionality, mathematical assistants can be running in the background to act as “copilots” and “mathematical grammar-checkers” — as the current generation of copilots [^6] Such as GitHub’s Copilot: https://github.com/features/copilot does successfully for programmers.
  8. 8.
    Mixed-mode interactions, where translation between formal and informal mathematics needs to be performed. These are particularly important for the combination of ITPs and general-purpose LLMs, which has recently picked up interest (e.g., (jiang2022draft; first2023baldur; frieder2023llms)).
@@
    Interactive, where the prompter (whose mathematical abilities can span a wide spectrum) solves mathematical questions together with the language model in an interactive way as in collins2023evaluating.
In practice, typical mathematical workflows require a mixture of different question and session types. Even if a question is not open-ended, the amount of mathematical detail and rigor the answer should contain will depend on the user’s background. In part, this should be deduced from the way the question is phrased. Currently, no datasets offer support for this. We note that with the exception of collins2023evaluating and the collection of CheckMate, almost all datasets are zero-shot. Interactive training modes are extensively employed in a formal language setup. IMProofBench evaluates research-level problems in an agentic setting with tools (web search and computer algebra), combining automated scoring on subproblems with expert grading of full proofs (improofbench2025).
@@
  • Refactoring formal mathematics to yield nicer-looking proof is a significant undertaking. A tool that takes as input a formal proof and refactors it along desired criteria would be a welcome addition to the toolbox of researchers working in formalizing mathematics. To our knowledge, the only work to date that attempts this is the ImProver (ahuja2024improver) for Lean, which allows formal proof creation of specific length.
@@ 3.1.4 Contamination
When releasing a dataset, control is lost over whether the dataset ends up as training data for machine learning models. For several of the state-of-the-art models, such as GPT-4 (openai2023gpt4), that perform well on benchmarks, no information is available in the ensuing on whether efforts were made to ensure that training datasets were decontaminated and no datapoints from the evaluation benchmark were included in the training set. For Gemini 1.5 (reid2024gemini), Qwen2 (yang2024qwen2) and DeepSeekMath (shao2024deepseekmath), only $n$-gram decontamination approaches have been tested. For complex datapoints, such as math, $n$-grams are unfortunately not always sufficient to ensure the training dataset is clean. There is evidence that suggests that several such modern models are already contaminated (xu2024benchmarking). While mitigating approaches have been proposed, such as generating new data, either from scratch (mishra2024mathcamps) or by using existing datasets as seeds (zhou2024your), it is not clear whether these can scale to more sophisticated mathematics, as the methods were tested using mathematics on the level of grade-school, and it is unclear how well they would scale to much higher levels of mathematical abstractions and problem-solving difficulty (the two main metrics of difficulty, as outlined in 3.1.1.
@@
If the dataset is sufficiently large, it is advisable to keep a part of the dataset hidden from the public so that the dataset creators can compare how well newly-released, publicly accessible open-weight models score in the public versus the hidden part of the dataset. Diverging scores on two splits of the dataset can indicate contamination. Care must be taken that datapoints from the hidden datasets are from the same distribution compared. For mathematics, this means at minimum: Same domain, difficulty range, and proof technique. This can be challenging: Given a specific problem $P$, it takes effort to source a new problem $P^{\star}$, that on all relevant metrics (including the mentioned ones) is similar to $P$, but at the same time is not completely analogous to $P$ – as it would be by, e.g., merely changing items in the problem statement that have a negligible on the proof (this is an approach followed by GSM8K-Symbolic dataset (mirzadeh2024gsm), which nonetheless turned out to be challenging for LLMs, highlighting their current limited reasoning capabilities). The approach advocated here has limitations for models released as-a-service (la2024language), via APIs or GUIs, as running the hidden dataset through risks it automatically being used at a later stage as training data for those models. RealMath proposes a continuously refreshed benchmark mined from arXiv papers and forum discussions, which reduces the half-life of static public test sets (realmath2025).
3.1.5 Binned Sankey diagrams
Math benchmarks are often reported with aggregate statistics such as accuracy, pass@$k$, or majority-vote accuracy. These statistics compress item-level outcomes into a small number of scalars. On mixed benchmarks, an average gain can coincide with regressions on specific slices (topic, modality, or answer format), which the average does not show. When assessing model improvement over time, this information conveys both too little detail and is also hard to read. This issue is not solely applicable to mathematics, but is very prevalent in this case.
Recommendation.
To support better visualizations, we recommend the use of binned Sankey diagrams. These are item-level transition plots between two evaluations. Each benchmark item is assigned to an outcome bin under a baseline evaluation (for example, model $M_{t}$) and to an outcome bin under a second evaluation (model $M_{t+1}$). The diagram draws flows from baseline bins to new bins, with band widths proportional to the number of items making each transition. Sankey-style migration plots have been used in other evaluation contexts; see Fig. 2 in frieder2023mathematical.
For mathematics, the visualization works best in the following way: Choose two evaluation points and define a small set of outcome bins. The minimal binning is binary (correct/incorrect). If annotations exist, bins can reflect error types (frieder2023mathematical) (forsexample, algebra error, invalid step, wrong final answer) or partial credit (for example, solved, nearly solved, off-track). If the evaluation involves tools, bins can separate solved without tools, solved with tools, and failed. The resulting diagram makes regressions explicit. A flow from solved to failed remains visible even when net accuracy increases. For heterogeneous benchmarks, the same transition plot can be computed within slices such as topic, difficulty level, or representation (symbolic, word problem, diagram-based, or formal proof assistant). Slice-level transitions show whether gains are concentrated in a subset of items and whether any slice exhibits systematic regressions.
Benchmark reports can include a binned Sankey diagram alongside aggregate metrics when comparing model versions or training regimes. This requires storing per-item outcomes (and, when applicable, slice labels) so that transitions can be reproduced. Using the same bins across releases supports longitudinal comparisons.